Three people pull simultaneously on a stubborn donkey. Jack pulls directly ahead of the donkey with a force of 81.9 N, Jill pulls with 69.1 N in a direction 45° to the left, and Jane pulls in a direction 45° to the right with 175 N. (Since the donkey is involved with such uncoordinated people, who can blame it for being stubborn?) Find the magnitude of the net force the people exert on the donkey.
Answer #1

We have shown the forces in the figure
F1 by Jack = 81.9 N
F2 by Jill = 69.1 N
F3 by Jane = 175 N
Now taking their component as shown in the figure.
Therefore net force in the X direction
FX = F1 + F2Cos45 + F3Cos45 = 81.9 + 69.1Cos45 + 175Cos45 = 254.505 N
Now in the Y direction
FY = F2Sin45 - F3Sin45 = (69.1 -175)Sin45 = - 74.88 N
Now the net force will be
F = \sqrt{F_{X}^{2} + F_{Y}^{2}}
F = 265.29 N
F, Sin45 2 5F Cos45 45 F- F2 Sin45